OpenStudy (anonymous):

\[\sum_{k=1}^{4}(k+7)^{2}\] what are the terms of the series?

OpenStudy (anonymous):

k=1 (k+7)^2=64 k=2 (k+7)^2=81 k=3 (k+7)^2=100 k=4 (k+7)^2=121 Total: 366 What do you mean by "series?"

OpenStudy (anonymous):

replace \(k\) successively by \(1,2,3,4\) and see what you get

OpenStudy (anonymous):

to add, stick a plus sign between them